# On the Closed-Form of Flow Matching: Generalization Does Not Arise from Target Stochasticity

### v1

@@ Abstract
Modern deep generative models can now produce high-quality synthetic samples that are often indistinguishable from real training data. A growing body of research aims to understand why recent methods – such as diffusion and flow matching techniques – generalize so effectively. Among the proposed explanations are the inductive biases of deep learning architectures and the stochastic nature of the conditional flow matching loss. In this work, we rule out the latter – the noisy nature of the loss – as a primary contributor to generalization in flow matching. First, we empirically show that in high-dimensional settings, the stochastic and closed-form versions of the flow matching loss yield nearly equivalent losses. Then, using state-of-the-art flow matching models on standard image datasets, we demonstrate that both variants achieve comparable statistical performance, with the surprising observation that using the closed-form can even improve performance.
@@ 1 Introduction
Recent deep generative models, such as diffusion (Sohl-Dickstein et al., 2015; Ho et al., 2020a; Song et al., 2021) and flow matching models (Lipman et al., 2023; Albergo and Vanden-Eijnden, 2023; Liu et al., 2023), have achieved remarkable success in synthesizing realistic data across a wide range of domains. State-of-the-art diffusion and flow matching methods are now capable of producing multi-modal outputs that are virtually indistinguishable from human-generated content, including images (Stability AI, 2023), audio (Borsos et al., 2023), video (Villegas et al., 2022; Brooks et al., 2024), and text (Gong et al., 2023; Xu et al., 2025).
A central question in deep generative modeling concerns the generalization capabilities and underlying mechanisms of these models. Generative models generalization remains a puzzling phenomenon, raising a number of challenging and unresolved questions: whether generative models truly generalize is still the subject of active debate. On one hand, several studies (Carlini et al., 2023; Somepalli et al., 2023b, a; Dar et al., 2023) have shown that large diffusion models are capable of memorizing individual samples from the training set, including licensed photographs, trademarked logos, and sensitive medical data.
On the other hand, Kadkhodaie et al. (2024) have empirically demonstrated that while memorization can occur in low-data regimes, diffusion models trained on a sufficiently large dataset exhibit clear signs of generalization. Taken together, recent work points to a sharp phase transition between memorization and generalization (Yoon et al., 2023; Zhang et al., 2024). Multiple theories have been proposed to explain the puzzling generalization of diffusion and flow matching models. On the one hand, Kadkhodaie et al. (2024); Kamb and Ganguli (2024); Ross et al. (2025) suggested a geometric framework to understand the inductive bias of modern deep convolutional networks on images. On the other hand, Vastola (2025) suggested that generalization is due to the noisy nature of the training loss. In this work, we clearly answer the following question:
@@
  • We challenge the prevailing belief that generalization in flow matching stems from an inherently noisy loss (Section 3.1). This assumption, largely supported by studies in low-dimensional settings, fails to hold in realistic high-dimensional data regimes.
  • Instead, we observe that generalization in flow matching emerges precisely when the limited-capacity neural network fails to approximate the exact velocity field (Section 3.2).
  • We identify two critical time intervals – at early and late times – where this approximation breaks down (Section 3.3). We show that generalization arises predominantly early along flow matching trajectories, aligning with the transition from the stochastic to the deterministic regime of the flow matching objective.
  • Finally, on standard image datasets (CIFAR-10 and CelebA), we show that explicitly regressing against the exact velocity field does not impair generalization and can, in some cases, enhance it (Section 4).
The manuscript is organized as follows. Section 2 reviews the fundamentals of conditional flow matching and recalls the closed-form of the “optimal” velocity field. Leveraging the closed-form expression of the flow matching velocity field, Section 3 investigates the key sources of generalization in flow matching. In Section 4, we introduce a learning algorithm based on the closed-form formula. Related work is discussed in detail in Section 5.
@@ 2 Recalls on conditional flow matching
Let $p_{0}$ be the source distribution and $p_{\mathrm{data}}$ the data distribution. We are given $n$ data points $x^{(1)},\dots,x^{(n)}\sim p_{\mathrm{data}}$, $x^{(i)}\in\mathbb{R}^{d}$. The goal of flow matching is to find a velocity field $u:\mathbb{R}^{d}\times[0,1]\to\mathbb{R}^{d}$, such that, if one solves on $[0,1]$ the ordinary differential equation
@@
then the law of $x(1)$ when $x_{0}\sim p_{0}$ is $p_{\mathrm{data}}$ (in which case we say that $u$ transports $p_{0}$ to $p_{\mathrm{data}}$). For every value of $t$ between 00 and $1$, the law of $x(t)$ defines a probability path, denoted $p(\cdot|t)$ that progressively goes from $p_{0}$ to $p_{\mathrm{data}}$. If one knows $u$, new samples can then be generated by sampling $x_{0}$ from $p_{0}$, solving the ordinary differential equation, and using $x(1)$ as the generated point.
In conditional flow matching, finding such a $u$ is achieved in the following way.
@@
    First, a conditioning variable $z$ independent of $t$ is defined, e.g. $z=x_{1}\sim p_{\mathrm{data}}$,
@@
    Then, a conditional probability path $p(\cdot|z,t)$ is chosen, e.g.555this choice is only valid when $p_{0}=\mathcal{N}(0,\operatorname*{\mathrm{Id}})$, which for simplicity we will assume in the sequel; more generic choices are possible and we refer the reader to Lipman et al. (2024); Gagneux et al. (2025) for deeper introductions to flow matching $p(\cdot|z=x_{1},t)=\mathcal{N}(tx_{1},(1-t)^{2}\cdot\operatorname*{\mathrm{Id}})$.
@@
Through the continuity equation (Lipman et al., 2024, Sec. 3.5), the choice of $p(\cdot|z,t)$ defines a conditional velocity field $u^{\mathrm{cond}}(x,z,t)$. In the setting of (i) and 5, we can for example choose $u^{\mathrm{cond}}(x,z=x_{1},t)=\frac{x_{1}-x}{1-t}$. By marginalization against $z$, conditional flow matching defines a probability path $p(\cdot|t)$ and a “total” velocity field
$u^{\star}(x,t)=\mathbb{E}_{z|x,t}\>u^{\mathrm{cond}}(x,z,t)\enspace,$ (2)
@@
which turns out to transport $p_{0}$ to $p_{\mathrm{data}}$ (Lipman et al., 2023, Thm. 1). Hence $u^{\star}$ could be approximated by a neural network $u_{\theta}:\mathbb{R}^{d}\times[0,1]\to\mathbb{R}^{d}$ with parameters $\theta$ by minimizing
@@
$\mathcal{L}_{\mathrm{FM}}(\theta)=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}([0,1])\\ x_{t}\sim p(\cdot|t)\end{subarray}}\|u_{\theta}(x_{t},t)-u^{\star}(x_{t},t)\|^{2}\enspace,$ (3)
@@
but the usual observation is that $u^{\star}$ is intractable. As a remedy, Lipman et al. (2023, Thm. 2) show that $\mathcal{L}_{\mathrm{FM}}(\theta)$ is equal up to a constant to the conditional flow matching loss, which for the choices (i) and 5 made above reads
@@
$\mathcal{L}_{\mathrm{CFM}}(\theta)=\mathbb{E}_{\begin{subarray}{c}x_{0}\sim p_{0}\\ x_{1}\sim p_{\mathrm{data}}\\ t\sim\mathcal{U}([0,1])\end{subarray}}\|u_{\theta}(x_{t},t)-u^{\mathrm{cond}}(x_{t},z=x_{1},t)\|^{2},\,\mathrm{where}\,x_{t}:=(1-t)x_{0}+tx_{1}\,.$ (4)
@@
Note that, since $x_{t}:=(1-t)x_{0}+tx_{1}$, we have that $u^{\mathrm{cond}}(x_{t},z=x_{1},t)$ is both equal to $\frac{x_{1}-x}{1-t}$and to $x_{1}-x_{0}$. The objective $\mathcal{L}_{\mathrm{CFM}}$ is easy to approximate, since it is easy to sample from $p_{0}$ and $\mathcal{U}([0,1])$; sampling from $p_{\mathrm{data}}$ is approximated by sampling from $\hat{p}_{\mathrm{data}}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{x^{(i)}}$. Although it seems natural, replacing $p_{\mathrm{data}}$ by $\hat{p}_{\mathrm{data}}$ in (4) has a very important consequence: it makes the minimizer $\hat{u}^{\star}$ of $\mathcal{L}_{\mathrm{FM}}$ available in closed-form, which we recall below.
Proposition 1 (Velocity Closed-form Formula).
@@
When $p_{\mathrm{data}}$ is replaced by $\hat{p}_{\mathrm{data}}$, with the previous choices, the velocity field in (2) has a closed-form formula:
$\hat{u}^{\star}(x,t)=\sum_{i=1}^{n}u^{\mathrm{cond}}(x,z=x^{(i)},t)\cdot\frac{p(x|z=x^{(i)},t)}{\sum_{i^{\prime}=1}^{n}p(x|z=x^{(i^{\prime})},t)}\enspace.$ (5)
@@
The notation $\hat{u}^{\star}$ emphasizes the velocity field is optimal for the empirical probability distribution $\hat{p}_{\mathrm{data}}$, not the true one $p_{\mathrm{data}}$. Since $u^{\mathrm{cond}}(x,z=x^{(i)},t)\propto x^{(i)}-x$, the optimal velocity field $\hat{u}^{\star}$ is a weighted average of the $n$ different directions $x^{(i)}-x$. More specifically, it writes as follows:
$\hat{u}^{\star}(x,t)=\sum_{i=1}^{n}\lambda_{i}(x,t)\frac{x^{(i)}-x}{1-t}\enspace,$ (6)
with $\lambda(x,t)=\mathrm{softmax}((-\frac{1}{2(1-t)^{2}}\|x-tx^{(i^{\prime})}\|^{2})_{i^{\prime}=1,\ldots,n})\in\mathbb{R}^{n}$. This formula also holds for a Gaussian source distribution and $z\sim p_{0}\times\hat{p}_{\mathrm{data}}$ (Section A.1).
This result can be found in various previous works, e.g., Kamb and Ganguli (2024, Eq. 3), Biroli et al. (2024), Gao and Li (2024), Li et al. (2024) or Scarvelis et al. (2025). Note that the closed-form formula (5) is also valid for other choices of continuous distribution $p_{0}$ (e.g., the uniform distribution).
@@
From Equation 6, as $t\to 1$, the velocity field $\hat{u}^{\star}$ diverges at any point $x$ that does not coincide with one of the training samples $x^{(i)}$, and it points in the direction of the nearest $x^{(i)}$. This creates a paradox: solving the ordinary differential equation (1) with the velocity field $\hat{u}^{\star}$ can only produce training samples $x^{(i)}$ (see Gao and Li 2024, Thm. 4.6 for a formal proof). Therefore, in practice, exactly minimizing the conditional flow matching loss would result in $u_{\theta}=\hat{u}^{\star}$, meaning the model memorizes the training data and fails to generalize. This naturally yields the following question:
@@ 3 Investigating the key sources of generalization
In this section, we investigate the key sources of flow matching generalization using the closed-form formula of its velocity field. First in Section 3.1 we challenge the claim that generalization stems from the stochastic approximation $u^{\mathrm{cond}}$ of the optimal velocity field $\hat{u}^{\star}$. Then, in Section 3.2 we show that generalization arises when $u_{\theta}$ fails to approximate the perfect velocity $\hat{u}^{\star}$. Interestingly, the target velocity estimation particularly fails at two critical time intervals. Section 3.3 shows that one of these critical times is particularly important for generalization.
@@ 3.1 Target stochasticity is not what you need
(a) Stochasticity vs. non-stochasticity
(b) Non stochasticity of $\hat{u}^{\star}$ for high-dimensional real data.
@@
Figure 1: We challenge the hypothesis that target stochasticity plays a major role in flow matching generalization. The histograms of the cosine similarities between $\hat{u}^{\star}((1-t)x_{0}+tx_{1},t)$ and $u^{\mathrm{cond}}((1-t)x_{0}+tx_{1},z=x_{1},t)=x_{1}-x_{0}$ are displayed for various time values $t$ and two datasets. For real, high-dimensional data, non-stochasticity arises very early (before $t=0.2$ for CIFAR-10 with dimension $(3,32,32)$).
@@
One recent hypothesis is that generalization arises from the fact that the regression target $u^{\mathrm{cond}}$ of conditional flow matching is only a stochastic estimate of $\hat{u}^{\star}$. The fact that the target regression objective only equals the true objective on average is referred to by Vastola (2025) as “generalization through variance”. To challenge this assumption, we leverage Proposition 1, which states that the optimal velocity field $\hat{u}^{\star}(x,t)$ is a weighted sum of the $n$ values of $u^{\mathrm{cond}}(x,t,z=x^{(i)})=\frac{x^{(i)}-x}{1-t}$, for $i\in[n]$, and show that, after a small time value $t$, this average is in practice equal to a single value in the expectation (see Figures 1(a) and 1(b)).
@@
Comments on Figure 1(b). To produce Figure 1(b), we sample $256$ pairs $(x_{0},x_{1})$ from $p_{0}\times\hat{p}_{\mathrm{data}}$. For each value of $t$, we compute the cosine similarity between the optimal velocity field $\hat{u}^{\star}((1-t)x_{0}+tx_{1},t)$ and the conditional target $u^{\mathrm{cond}}((1-t)x_{0}+tx_{1},z=x_{1},t)=x_{1}-x_{0}$. The resulting similarities are aggregated and shown as histograms. The top row displays the results for the two-moons toy dataset ($d=2$), and the bottom row displays the results for the CIFAR-10 dataset (Krizhevsky and Hinton 2009, $d=3072$); $n=50$k for both. As $t$ increases, the histograms become increasingly concentrated around $1$, indicating that $\hat{u}^{\star}$ aligns closely with a single conditional vector $u^{\mathrm{cond}}$. From Equation 6, this corresponds to a collapse towards 0 of all but one of the softmax weights $\lambda_{i}(x_{t},t)$. This time corresponds to the collapse time studied by Biroli et al. (2024) for diffusion; we discuss the connection in the related works (Section 5). On the two-moons toy dataset, this transition occurs for intermediate-to-large values of $t$, echoing the observations made in low-dimensional settings by Vastola (2025, Figure 1). In contrast, for high-dimensional real datasets, $\hat{u}^{\star}(x,t)$ aligns with a single conditional velocity field $x^{(i)}-x$, even at early time steps, suggesting that the non-stochastic regime dominates most of the generative process. This key difference between low- and high-dimensional data suggests that the transition time between the stochastic and non-stochastic regimes is strongly influenced by the dimensionality of the data.
@@
Dimension dependence. To further illustrate the strong impact of dimensionality, Figure 2 reports the proportion of samples $x_{t}$ (from a batch of 256) for which the cosine similarity between $\hat{u}^{\star}$ and $u^{\mathrm{cond}}\propto x^{(i)}-x$ exceeds $0.9$, as a function of time $t$. This analysis is performed across multiple spatial resolutions of the Imagenette dataset, obtaining $d\times d$ images by spatial subsampling.
Figure 2 reveals a sharp transition: as the dimensionality increases, the proportion of high-cosine matches rapidly converges to 100%.
Figure 2: Alignment of $\hat{u}^{\star}$ and $u^{\mathrm{cond}}$ over time for varying image dimensions $d$ on Imagenette.
A practical implication of this behavior is that, for sufficiently large $t$, if $x_{0}\sim p_{0}$ and $x^{(i)}\sim\hat{p}_{\mathrm{data}}$, then $\hat{u}^{\star}((1-t)x_{0}+tx^{(i)},t)$ is approximately proportional to $x^{(i)}-x$. Consequently, regressing on $x^{(i)}$ or on the conditional velocity $x_{1}-x_{0}$ becomes effectively equivalent. Section 4 investigates how to learn regressing against optimal velocity field $\hat{u}^{\star}$, and empirically shows similar results between stochastic and non-stochastic targets.

### v2

@@ Abstract
Modern deep generative models can now produce high-quality synthetic samples that are often indistinguishable from real training data. A growing body of research aims to understand why recent methods, such as diffusion and flow matching techniques, generalize so effectively. Among the proposed explanations are the inductive biases of deep learning architectures and the stochastic nature of the conditional flow matching loss. In this work, we rule out the noisy nature of the loss as a key factor driving generalization in flow matching. First, we empirically show that in high-dimensional settings, the stochastic and closed-form versions of the flow matching loss yield nearly equivalent losses. Then, using state-of-the-art flow matching models on standard image datasets, we demonstrate that both variants achieve comparable statistical performance, with the surprising observation that using the closed-form can even improve performance.
@@ 1 Introduction
Recent deep generative models, such as diffusion (sohlDickstein2015diffusion; Ho2020; song2020diffusion) and flow matching models (lipman2023flow; albergo2023stochasticinterpolant; liu2023rectifiedflow), have achieved remarkable success in synthesizing realistic data across a wide range of domains. State-of-the-art diffusion and flow matching methods are now capable of producing multi-modal outputs that are virtually indistinguishable from human-generated content, including images (StablediffXL), audio (Borsos2023), video (Villegas2022; videoworldsimulators2024), and text (Gong2022diffusiontext; Xu2025energytext).
A central question in deep generative modeling concerns the generalization capabilities and underlying mechanisms of these models. Generative models generalization remains a puzzling phenomenon, raising a number of challenging and unresolved questions: whether generative models truly generalize is still the subject of active debate. On one hand, several studies (Carlini2023; somepalli2023diffusion; Somepalli2023understanding; Dar2023investigating) have shown that large diffusion models are capable of memorizing individual samples from the training set, including licensed photographs, trademarked logos, and sensitive medical data.
On the other hand, kadkhodaie2024generalization have empirically demonstrated that while memorization can occur in low-data regimes, diffusion models trained on a sufficiently large dataset exhibit clear signs of generalization. Taken together, recent work points to a sharp phase transition between memorization and generalization (yoon2023diffusion; zhang2024emergencerepro). Multiple theories have been proposed to explain the puzzling generalization of diffusion and flow matching models. On the one hand, kadkhodaie2024generalization; kamb2024analytic; ross2025memorization suggested a geometric framework to understand the inductive bias of modern deep convolutional networks on images. On the other hand, Vastola2025generalization suggested that generalization is due to the noisy nature of the training loss. In this work, we clearly answer the following question:
@@
  • We challenge the prevailing belief that generalization in flow matching stems from an inherently noisy loss (Section˜3.1). This assumption, largely supported by studies in low-dimensional settings, fails to hold in realistic high-dimensional data regimes.
  • Instead, we observe that generalization in flow matching emerges precisely when the limited-capacity neural network fails to approximate the optimal closed-form velocity field (Section˜3.2).
  • We identify two critical time intervals, at early and late times, where neural networks fail to approximate the optimal velocity field (Section˜3.3). We show that generalization arises predominantly early along flow matching trajectories, aligning with the transition from the stochastic to the deterministic regime of the flow matching objective.
  • Finally, on standard image datasets (CIFAR-10 and CelebA), we show that explicitly regressing against the optimal closed-form velocity field does not impair generalization and can, in some cases, enhance it (Section˜4).
The manuscript is organized as follows. Section˜2 reviews the fundamentals of conditional flow matching and recalls the closed-form of the “optimal” velocity field. Leveraging the closed-form expression of the flow matching velocity field, Section˜3 investigates the key sources of generalization in flow matching. In Section˜4, we introduce a learning algorithm based on the closed-form formula. Related work is discussed in detail in Section˜5.
@@ 2 Recalls on conditional flow matching
Let $p_{0}=\mathcal{N}(0,\operatorname*{\mathrm{Id}})$ be the source distribution111the choice $p_{0}=\mathcal{N}(0,\operatorname*{\mathrm{Id}})$ is made for simplicity; more generic choices are possible and the reader can refer to Lipman2024guide; Gagneux2025visual_flow_matching; diffusion_fm_blogpost for deeper introductions to flow matching. and $p_{\mathrm{data}}$ the data distribution. We are given $n$ data points $x^{(1)},\dots,x^{(n)}\sim p_{\mathrm{data}}$, $x^{(i)}\in\mathbb{R}^{d}$. The goal of flow matching is to find a velocity field $u:\mathbb{R}^{d}\times[0,1]\to\mathbb{R}^{d}$, such that, if one solves on $[0,1]$ the ordinary differential equation
@@
then the law of $x(1)$ when $x_{0}\sim p_{0}$ is $p_{\mathrm{data}}$: one says that $u$ transports $p_{0}$ to $p_{\mathrm{data}}$. For every value of $t$ between 0 and $1$, the law of $x(t)$ defines a probability path, denoted $p(\cdot|t)$ that progressively transforms $p_{0}$ to $p_{\mathrm{data}}$. If one knows the velocity field $u$, new samples can then be generated by sampling $x_{0}$ from $p_{0}$, solving the ordinary differential equation, and using $x(1)$ as the generated point.
In conditional flow matching, finding such a velocity field $u$ is achieved in the following way.
@@
    First, define a conditioning variable $z$ independent of $t$, e.g., $z=x_{1}\sim p_{\mathrm{data}}$,
@@
    Then, chose a conditional probability path $p(\cdot|z,t)$, e.g., $p(\cdot|z=x_{1},t)=\mathcal{N}(tx_{1},(1-t)^{2}\operatorname*{\mathrm{Id}})$.
@@
Through the continuity equation (Lipman2024guide, Sec. 3.5), the choice (ii) of the conditional probability path $p(\cdot|z,t)$ defines a conditional velocity field $u^{\mathrm{cond}}(x,z,t)$. With the choices (i) and (ii), the conditional velocity field writes
$u^{\mathrm{cond}}(x,z=x_{1},t)=\frac{x_{1}-x}{1-t}\kern 5.0pt.$ (2)
The choice (ii) of the conditional probability paths $p(\cdot|z=x_{1},t)$ fully defines a probability path $p(\cdot|t)$ (by marginalization against $z$) and thus defines an optimal velocity field $u^{\star}$ (through the continuity equation), that transports $p_{0}$ to $p_{\mathrm{data}}$ (lipman2023flow, Thm. 1)
$u^{\star}(x,t)=\mathbb{E}_{z|x,t}\>u^{\mathrm{cond}}(x,z,t)\kern 5.0pt.$ (3)
@@
Hence, the optimal velocity $u^{\star}$ could be approximated by a neural network $u_{\theta}:\mathbb{R}^{d}\times[0,1]\to\mathbb{R}^{d}$ with parameters $\theta$ by minimizing
@@
$\mathcal{L}_{\mathrm{FM}}(\theta)=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}([0,1])\\ x_{t}\sim p(\cdot|t)\end{subarray}}\|u_{\theta}(x_{t},t)-u^{\star}(x_{t},t)\|^{2}\kern 5.0pt.$ (4)
@@
However, $u^{\star}$ is usually (believed) intractable, as a remedy, lipman2023flow showed that $\mathcal{L}_{\mathrm{FM}}(\theta)$ is equal, up to a constant, to the conditional flow matching loss. With the choices (i) and (ii) made above, the conditional flow matching loss reads
@@
$\mathcal{L}_{\mathrm{CFM}}(\theta)=\mathbb{E}_{\begin{subarray}{c}x_{0}\sim p_{0}\\ x_{1}\sim p_{\mathrm{data}}\\ t\sim\mathcal{U}([0,1])\\ \end{subarray}}\|u_{\theta}(x_{t},t)-\underbrace{u^{\mathrm{cond}}(x_{t},z=x_{1},t)}_{=\frac{x_{1}-x_{t}}{1-t}=x_{1}-x_{0}}\|^{2},$ (5)
@@
where $x_{t}:=(1-t)x_{0}+tx_{1}$. The objective $\mathcal{L}_{\mathrm{CFM}}$ is easy to approximate, since it is easy to sample from $p_{0}=\mathcal{N}(0,\operatorname*{\mathrm{Id}})$ and $\mathcal{U}([0,1])$; sampling from $p_{\mathrm{data}}$ is approximated by sampling from $\hat{p}_{\mathrm{data}}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{x^{(i)}}$. Although it seems natural, replacing $p_{\mathrm{data}}$ by $\hat{p}_{\mathrm{data}}$ in (5) has a very important consequence: it makes the minimizer $\hat{u}^{\star}$ of $\mathcal{L}_{\mathrm{FM}}$ available in closed-form, which we recall below.
Proposition 1 (Closed-form Formula of the Optimal Velocity).
@@
When $p_{\mathrm{data}}$ is replaced by $\hat{p}_{\mathrm{data}}$, with the previous choices (i) and (ii), the optimal velocity field $\hat{u}^{\star}$ in (3) has a closed-form formula:
$\hat{u}^{\star}(x,t)=\sum_{i=1}^{n}\lambda_{i}(x,t)\frac{x^{(i)}-x}{1-t}\kern 5.0pt,$ (6)
with $\lambda(x,t)=\mathrm{softmax}((-\frac{\|x-tx^{(j)}\|^{2}}{2(1-t)^{2}})_{j=1,\ldots,n})\in\mathbb{R}^{n}.$
@@
The notation $\hat{u}^{\star}$ emphasizes the velocity field is optimal for the empirical probability distribution $\hat{p}_{\mathrm{data}}$, not the true one $p_{\mathrm{data}}$. Since $u^{\mathrm{cond}}(x,z=x^{(i)},t)\propto x^{(i)}-x$, the optimal velocity field $\hat{u}^{\star}$ is a weighted average of the $n$ different directions $x^{(i)}-x$. Note that the closed-form formula in Equation˜6 can be found in various previous works, e.g., kamb2024analytic, Biroli2024, gao2024flow, li2024goodscore or Scarvelis2023, and can be generalized to other choices of continuous distribution $p_{0}$ (e.g., the uniform distribution, see Section˜A.1).
@@
From Equation˜6, as $t\to 1$, the velocity field $\hat{u}^{\star}$ diverges at any point $x$ that does not coincide with one of the training samples $x^{(i)}$, and it points in the direction of the nearest $x^{(i)}$. This creates a paradox: solving the ordinary differential equation (1) with the velocity field $\hat{u}^{\star}$ can only produce training samples $x^{(i)}$ (see gao2024flow for a formal proof). Therefore, in practice, exactly minimizing the conditional flow matching loss would result in $u_{\theta}=\hat{u}^{\star}$, meaning the model memorizes the training data and fails to generalize. This naturally yields the following question:
@@ 3 Investigating the key sources of generalization
In this section, we investigate the key sources of flow matching generalization using the closed-form formula of its velocity field. First in Section˜3.1 we challenge the claim that generalization stems from the stochastic approximation $u^{\mathrm{cond}}$ of the optimal velocity field $\hat{u}^{\star}$. Then, in Section˜3.2 we show that generalization arises when $u_{\theta}$ fails to approximate the perfect velocity $\hat{u}^{\star}$. Interestingly, the target velocity estimation particularly fails at two critical time intervals. Section˜3.3 shows that one of these critical times is particularly important for generalization.
@@ 3.1 Target stochasticity is not what you need
(a) Non stochasticity of $\hat{u}^{\star}$ for high-dimensional real data.
(b) Stochasticity vs. non-stochasticity
(c) Dimension Dependence
@@
Figure 1: We challenge the hypothesis that target stochasticity plays a major role in flow matching generalization. In Figure˜1(a), the histograms of the cosine similarities between $\hat{u}^{\star}((1-t)x_{0}+tx_{1},t)$ and $u^{\mathrm{cond}}((1-t)x_{0}+tx_{1},z=x_{1},t)=x_{1}-x_{0}$ are displayed for various time values $t$ and two datasets. For real, high-dimensional data, non-stochasticity arises very early (before $t=0.2$ for CIFAR-10 with dimension $(3,32,32)$). Figure˜1(c) displays the alignment between $\hat{u}^{\star}$ and $u^{\mathrm{cond}}$ over time for varying image dimensions $d$ on Imagenette.
@@
One recent hypothesis is that generalization arises from the fact that the regression target $u^{\mathrm{cond}}$ of conditional flow matching is only a stochastic estimate of $\hat{u}^{\star}$. The fact that the target regression objective only equals the true objective on average is referred to by Vastola2025generalization as “generalization through variance”. To challenge this assumption, we leverage Proposition˜1, which states that the optimal velocity field $\hat{u}^{\star}(x,t)$ is a weighted sum of the $n$ values of $u^{\mathrm{cond}}(x,t,z=x^{(i)})=\frac{x^{(i)}-x}{1-t}$, for $i\in[n]$, and show that, after a small time value $t$, this average is in practice equal to a single value in the expectation (see Figures˜1(b) and 1(a)).
@@
Comments on Figure˜1(a). To produce Figure˜1(a), we sample $256$ pairs $(x_{0},x_{1})$ from $p_{0}\times\hat{p}_{\mathrm{data}}$. For each value of $t$, we compute the cosine similarity between the optimal velocity field $\hat{u}^{\star}((1-t)x_{0}+tx_{1},t)$ and the conditional target $u^{\mathrm{cond}}((1-t)x_{0}+tx_{1},z=x_{1},t)=x_{1}-x_{0}$. The resulting similarities are aggregated and shown as histograms. The top row displays the results for the two-moons toy dataset ($d=2$), and the bottom row displays the results for the CIFAR-10 dataset (Krizhevsky2009, $d=3072$); n=50n=50k for both. As $t$ increases, the histograms become increasingly concentrated around $1$, indicating that $\hat{u}^{\star}$ aligns closely with a single conditional vector $u^{\mathrm{cond}}$. From Equation˜6, this corresponds to a collapse towards 0 of all but one of the softmax weights $\lambda_{i}(x_{t},t)$. This time corresponds to the collapse time studied by Biroli2024 for diffusion; we discuss the connection in the related works (Section˜5). On the two-moons toy dataset, this transition occurs for intermediate-to-large values of $t$, echoing the observations made in low-dimensional settings by Vastola2025generalization. In contrast, for high-dimensional real datasets, $\hat{u}^{\star}(x,t)$ aligns with a single conditional velocity field $x^{(i)}-x$, even at early time steps, suggesting that the non-stochastic regime dominates most of the generative process. This key difference between low- and high-dimensional data suggests that the transition time between the stochastic and non-stochastic regimes is strongly influenced by the dimensionality of the data.
@@
Comments on Figure˜1(c). To further illustrate the strong impact of dimensionality, Figure˜1(c) reports the proportion of samples $x_{t}$ (from a batch of 256) for which the cosine similarity between $\hat{u}^{\star}$ and $u^{\mathrm{cond}}\propto x^{(i)}-x$ exceeds $0.9$, as a function of time $t$. This analysis is performed across multiple spatial resolutions of the Imagenette dataset (Howard_Imagenette_2019), obtaining $\mathrm{dim}\times\mathrm{dim}$ images by spatial subsampling. Figure˜1(c) reveals a sharp transition: as the dimensionality increases, the proportion of high-cosine matches rapidly converges to 100%. A practical implication of this behavior is that, for sufficiently large $t$, if $x_{0}\sim p_{0}$ and $x^{(i)}\sim\hat{p}_{\mathrm{data}}$, then $\hat{u}^{\star}((1-t)x_{0}+tx^{(i)},t)$ is approximately proportional to $x^{(i)}-x$. Consequently, regressing on $x^{(i)}$ or on the conditional velocity $x_{1}-x_{0}$ becomes effectively equivalent. Section˜4 investigates how to learn regressing against optimal velocity field $\hat{u}^{\star}$, and empirically shows similar results between stochastic and non-stochastic targets.